Shift report 20030210 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Feb 10 2003 - 00:17:50 EST

  • Next message: Apache: "Shift report 20030210 08:00-16:00"
    ------------------------------------------------------------
    Date: Mon Feb 10 00:17:49 2003, Shift: 16:00-24:00
    Supervisor: YKL, with: MG
    Log book#14 pages 7 - 8
    
    Shift Summary:
    ------------------------------------------------------------
    
    1.Had a useful beam only in the last 3 hours.
    
    2 Data runs are 7853,7854, 7855(still running) with a total trigger 3 (MRS) of 64 k and trigger 6 (FS) of 26 k. 3. In the current run 7856 trigger 3 scale is set to 1. 4. IMPORTANT: Due to an error in the last shift, FFS has been set in a wrong way (8 deg with BFS at 12 deg). FOR THIS REASON THE RUNS 7845,7846,7847,7848 ARE UNUSEFUL AND HAVE TO BE REDONE. 
    
    3. T3 anode A3-yv tripped down several times. With help of PS the problem has been solved. Sum instructions how to solve this problem have been printed out.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Feb 10 2003 - 00:21:47 EST